The Ks of strontium sulfite, SrSO3, is 4.0 x 10^−8 at 25°C. If 1.0000 g of SrSO3 is stirred in 5.0 L of water until the solution is saturated and then filtered, What mass of SrSO3 would remain?
1
Expert's answer
2020-06-05T05:14:08-0400

Molar mass = 183.68 gms

Concentration of salt="c=1\/(183.68*5)= 1.088*10^{-3}M"

"K_{sp} = s^2=4*10^{-8}M \implies s= *10^{-4}M = 0.2*10^{-3}M"

Thus, concentration of SrSO3 remaining is "c-s= 0.888*10^{-3}M"

Moles remaining = "0.888*10^{-3}*5=4.444*10^{-3} mol"

Thus, amount of SrSO3 remaining = "4.444*10^{-3}*183.68=0.8163 gms."